Fermi National Laboratory
 

Accelerator Update for April 23, 2007 - April 25, 2007

 

Monday April 23

The day shift began with Operations monitoring the Tevatron (TeV) store 5369 (42.13E30) and the Recycler stash (292.4E10), with the Antiproton Source (Pbar) stacking (9.4E10), with the Linac, Booster, and Main Injector (MI) operating normally, with NuMI, MiniBooNE, and the Meson MTest experiments taking beam, and with D-Zero (D0) and the Collider Detector at Fermilab (CDF) using the colliding beams.

A Pbar expert began a study at 9:44 AM that halted stacking for about 30 minutes.

At 10:36 AM, a Linac RF station (LRF2) tripped off. It reset.

A Pbar and a MI expert began study periods at 12:37 PM.

The Pbar study ended at 2:04 PM.

The MI study ended at 3:16 PM.

At 3:56 PM, LRF3 tripped off. It reset.

Operations reported at 3:58 PM that the Recycler had stashed antiprotons twice during the day shift.

The MTest experiment stopped taking beam at 4:58 PM.

Operations terminated store 5369 at 5:05 PM.

Operations put the TeV into shot setup at 5:23 PM.

At 6:45 PM, Operations reported a LCW leak in the Pbar 95 LCW system. Operators and FESS personnel inspected CUB and the Pbar service buildings; they found no leak.

Operations established store 5370 at 7:39 PM with an initial luminosity of 276.4E30.

At 7:54 PM, Operations reported that humidity sensors in the upstream area of Pbar transport had registered an increase since the LCW leak had been noticed. Operations called in LCW technicians and prepared the area for an access.

At 8:05 PM, Operations stashed all the Pbar antiprotons into the Recycler in preparation for the access.

At 9:23 PM, Operations reported that the LCW techs found and fixed a hose leak on a quadrupole magnet (IQ19).

Operations had established beam to Pbar and NuMI by 9:48 PM.

At 10:39 PM, Operations reported that LRF3 had tripped a couple of times. It reset.

At 11:14 PM, a Linac Klystron RF station (KRF3) tripped off. It reset.

At 11:55 PM, Operations reported a P1 transfer line vacuum burst. Everything reset.

Operations reported at midnight that the Recycler had stashed twice during the evening shift.

Tuesday April 24

The midnight shift began with Operations monitoring store 5370 (139.27E30) and the stash (65.0E10), with Pbar stacking (39.7E10), and with NuMI and MiniBooNE taking beam.

At 12:46 AM, KRF6 tripped off. It reset.

Operations established beam to MTest at 6:15 AM.

Operations reported at 8:07 AM that the Recycler had stashed four times during the midnight shift.

At 9:22 AM, LRF3 tripped off. It reset.

An RF expert reported at 10:06 AM that the Accumulator RF station, ARF4, was repaired. It will be brought online on Wednesday.

At 12:03 PM, Pbar began an eight hour period of lattice studies. Both the Recycler and MI also conducted studies.

At 12:06 PM, LRF3 tripped off. Experts investigated and reset the station.

At 12:27 PM, a Safety Officer reported a sinkhole near the Pbar transfer line AP0 service building. Operations reported the problem to Work Central. FESS personnel reported that the hole was due to a broken sump discharge, which was going to be repaired at a later date. Safety made sure the hole was filled in.

At 12:48 PM, KRF4 tripped off. Linac techs changed the card for KRF4 in its VME crate. Concurrently, NuMI and MI experts set up for 11- batch slipstacked beam.

At 1:34 PM, KRF4 tripped off again. The card replacement didn’t fix the problem so the experts reinstalled the original card.

At 2:20 PM, LRF3 tripped off. It reset.

At 3:23 PM, MI tripped off due to vacuum burst, followed by a vacuum cascade in the MI-40 line. Everything reset and MI resumed accelerating beam at 3:31 PM.

Operations reported at 3:57 PM that the Recycler had stashed once during the day shift.

MI and NuMI resumed normal 7-batch operations at 4:41 PM.

Operation reported at 5:21 PM the following studies’ status: Pbar experts were using reverse protons for a tune-up, MI experts were conducting barrier bucket studies, and Recycler experts were setting up stochastic cooling so they could resume normal machine operations.

MTest reported at 6:02 PM that they were through with the beam. This concluded their run. Operations disabled the MTest beam.

At 8:15 PM, a TeV sector A49 separator’s output stopped. There were no sparks. Operators switched the device to local control and set its output by hand. Later, operators traced the problem to a shorted capacitor in its 15-volt power supply regulator. Operations called in an expert to make repairs.

Operations reported that the Pbar study had ended and stacking resumed around 8:45 PM.

Operations terminated store 5370 at 11:28 PM.

Experts replaced two TeV separator power supply regulator cards at 11:56 PM for A49H and A49V.

Wednesday April 25

The midnight shift began with Operations monitoring the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

Operations put the TeV into a dry squeeze at 12:15 AM.

Operations put the TeV into a 4x0 wet squeeze at 1:16 AM.

At 2:20 AM, LRF3 tripped off. It reset.

At 2:21 AM, KRF4 tripped off. It reset.

Operations put the TeV through another dry squeeze at 2:33 AM.

Operations put the TeV into shot setup at 3:06 AM.

At 4:54 AM, a MI RF station (MIRF9) tripped off and caused a bad beam position that tripped off the NuMI beam permit and also caused a Lambertson (LAM61) breaker to trip. Operations called in an expert to reset the breaker.

Operations established store 5374 at 4:59 AM with an initial luminosity of 241.7E30.

At 5:35 AM, Operations reported that MIRF9 had been bypassed.

NuMI resumed taking beam at 5:46 AM, after an expert reset the breaker for LAM61.

Operations reported at 8:10 AM that the Recycler had stashed once during the midnight shift.

The Plans for Wednesday

The plans for today are to repair MIRF9 and then resume stacking, stashing, and storing.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 04/25/2007    email Fermilab